# Regularized KL-Divergence for Well-Defined Function-Space Variational Inference in Bayesian Neural Networks

### v2

@@ Abstract
Bayesian neural networks (BNN) promise to combine the predictive performance of neural networks with principled uncertainty modeling important for safety-critical systems and decision making. However, posterior uncertainty estimates depend on the choice of prior, and finding informative priors in weight-space has proven difficult. This has motivated variational inference (VI) methods that pose priors directly on the function generated by the BNN rather than on weights. In this paper, we address a fundamental issue with such function-space VI approaches pointed out by Burt et al., (2020), who showed that the objective function (ELBO) is negative infinite for most priors of interest. Our solution builds on generalized VI (Knoblauch et al.,, 2019) with the regularized KL divergence (Quang,, 2019) and is, to the best of our knowledge, the first well-defined variational objective for function-space inference in BNNs with Gaussian process (GP) priors. Experiments show that our method incorporates the properties specified by the GP prior on synthetic and small real-world data sets, and provides competitive uncertainty estimates for regression, classification and out-of-distribution detection compared to BNN baselines with both function and weight-space priors.
@@ 1 Introduction
Neural networks have shown impressive results in many fields but fail to provide well-calibrated uncertainty estimates, which are essential in applications associated with risk, such as healthcare (Kompa et al.,, 2021; Abdullah et al.,, 2022) or finance (Bew et al.,, 2019; Wong,, 2023). Bayesian neural networks (BNNs) offer to combine the scalability and predictive performance of neural networks with principled uncertainty modeling by explicitly capturing epistemic uncertainty, i.e., uncertainty resulting from finite training data. While the choice of prior in the Bayesian framework strongly affects the uncertainty estimates later obtained from the posterior, specifying informative priors on BNN weights has proven difficult and is hypothesized to have limited their practical applicability (Knoblauch et al.,, 2019; Cinquin et al.,, 2021; Tran et al.,, 2022). For instance, the default isotropic Gaussian prior, which is often chosen for tractability rather than for the beliefs it carries (Knoblauch et al.,, 2019), is known to have pathological behavior in some cases (Tran et al.,, 2022). A promising approach to solve this issue is to place priors directly on the function represented by the BNN instead of the weights. While being technically more challenging, function-space priors allow incorporating interpretable knowledge into the inference, for instance enabling the use of the extensive Gaussian Process (GP) literature to improve prior selection and design (Williams and Rasmussen,, 2006).
@@
A recent line of work has focused on using function-space priors in BNNs with variational inference (VI) (Sun et al.,, 2019). VI is an appealing method because of its successful application to BNNs, its flexibility in terms of approximate posterior parameterization, and its scalability to large datasets and models (Hoffman et al.,, 2013; Blundell et al.,, 2015; Tomczak et al.,, 2020). Unfortunately, in the context of BNNs with function-space priors, the Kullbach-Leibler (KL) divergence term in the VI objective (ELBO) involves two intractabilities: (i) a supremum over infinitely many subsets and (ii) access to the density of the distribution of function represented by the BNN, which has no closed-form expression. Sun et al., (2019) propose to address problem (i) by approximating the supremum in the KL divergence by an expectation, and problem (ii) by using implicit score function estimators (which make this method difficult to use in practice (Ma and Hernández-Lobato,, 2021)). However, the problem is actually more severe. Not only is the KL divergence intractable, it is infinite in many cases of interest (Burt et al.,, 2020), such as when the prior is a non-degenerate GP or a BNN with a different architecture. Thus, in these (and many more) situations, the KL divergence cannot even be approximated. As a consequence, more recent work abandons using BNNs and instead uses deterministic neural networks to parameterize basis functions (Ma and Hernández-Lobato,, 2021) or a GP mean (Wild et al.,, 2022). The only prior work (Rudner et al., 2022b, ) on function-space VI in BNNs that overcomes the issue pointed out by Burt et al., (2020) does so by deliberately limiting itself to cases where the KL divergence is known to be finite (by defining the prior as the pushforward of a weight-space distribution). Therefore, the method by Rudner et al., 2022b suffers from the same issues regarding prior specification as any other weight-space inference method for BNNs.
@@
In this paper, we address the argument by Burt et al., (2020) that VI does not provide a valid objective for inference in BNNs with genuine function-space priors, and we propose to apply the framework of generalized variational inference (Knoblauch et al.,, 2019). We present a simple method for function-space inference with GP priors that builds on the so-called regularized KL divergence (Quang,, 2019), which generalizes the conventional KL divergence and is finite for any pair of Gaussian measures. We obtain a Gaussian measure for the variational posterior by considering the linearized BNN from Rudner et al., 2022b , and we are free to choose a prior from a large set of GPs which have an associated Gaussian measure on the considered function space. While the regularized KL divergence is still intractable, it can be consistently estimated from samples with a known error bound. We find that our method effectively incorporates the beliefs specified by GP priors into the inference process (see figure 1) and provides competitive performance compared to BNN baselines. To the best of our knowledge, our method is the first to provide a well-defined objective for function-space inference in BNNs with informative GP priors. Our contributions are summarized below:
@@
    We propose to use generalized VI with the so-called regularized KL divergence to mitigate the issue of an infinite KL divergence when using VI in BNNs with function-space priors.
@@
    We present a new, well-defined objective for function-space inference in the linearized BNN with GP priors, which results in a simple algorithm.
@@
    We analyze our method empirically on synthetic and real-world data sets, and find that it accurately captures structural properties specified by the GP prior and provides competitive uncertainty estimates for regression, classification, and out-of-distribution detection compared to BNN baselines with both function-space and weight-space priors.
The paper is structured as follows. section 2 introduces function-space VI in BNNs and the regularized KL divergence. section 3 presents our proposed method for generalized function-space VI (GFSVI) in BNNs. section 4 presents experimental results. We discuss related work in section 5 and conclude in section 6.
Figure 1: BNN inference on synthetic data (gray circles) with Gaussian process priors encoding different properties such as smoothness (increasing from Matern-1/2 to RBF) and periodicity (periodic).
@@ 2 Background
In this section, we provide background on function-space variational inference (VI) in BNNs and discuss the fundamental issue of an infinite KL-divergence. We then introduce the regularized KL divergence, which is the basis for our solution presented in Section 3.
@@ 2.1 Function-space VI in BNNs
We consider a neural network $f({\,\cdot\,;{\bm{w}}})$ with weights ${\bm{w}}\in^{p}$, and a data set $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ with features ${\bm{x}}_{i}\in\mathcal{X}\subset^{d}$ and values $y_{i}\in\mathcal{Y}$. Bayesian Neural Networks (BNNs) are specified further by a likelihood function $p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}=\prod_{i=1}^{N}p\mathopen{}\left(y_{i}\,|\,f({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}$ and—traditionally—a prior $p\mathopen{}\left({\bm{w}}\right)\mathclose{}$ on the weights, and one seeks the posterior distribution $p\mathopen{}\left({\bm{w}}\,|\,\mathcal{D}\right)\mathclose{}\propto p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}\,p\mathopen{}\left({\bm{w}}\right)\mathclose{}$. The method proposed in this paper builds on variational inference (VI), which approximates $p\mathopen{}\left({\bm{w}}\,|\,\mathcal{D}\right)\mathclose{}$ with a variational distribution $q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}$, whose variational parameters $\phi$ maximize the evidence lower bound (ELBO),
@@
where $D_{\text{KL}}$ is the Kullback-Leibler (KL) divergence
@@
Function-space variational inference
Since weights of neural networks are not interpretable, we abandon the weight-space prior $p({\bm{w}})$ and instead pose a prior $\mathbb{P}$ directly on the function $f\mathopen{}\left(\,\cdot\,;{\bm{w}}\right)\mathclose{}$, which we denote simply as $f$ when there is no ambiguity. Here, the symbol $\mathbb{P}$ denotes a probability measure that does not admit a density since the function space is infinite-dimensional. We compute the expected log-likelihood as in the first term of Eq. 2.1, but we calculate the KL-term (Eq. 2.2) using the pushforward of $q_{\phi}({\bm{w}})$ along the mapping ${\bm{w}}\mapsto f\mathopen{}\left(\,\cdot\,;{\bm{w}}\right)\mathclose{}$, which defines the variational measure $\mathbb{Q}_{\phi}$, resulting in the ELBO in function space,
@@
with $D_{\text{KL}}$ the KL divergence between measures given by
@@
Here, the Raydon-Nikodym derivative $d\mathbb{Q}_{\phi}/d\mathbb{P}$ generalizes the density ratio from Eq. 2.2. Like Eq. 2.1, the ELBO in Eq. 2.3 is a lower bound on the evidence (Burt et al.,, 2020). In fact, if $\mathbb{P}$ is the push-forward of $p({\bm{w}})$ then Eq. 2.3 is a tighter bound than Eq. 2.1 by the data processing inequality, $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}\leq D_{\mathrm{KL}}\mathopen{}\left(q_{\phi}\,\|\,p\right)\mathclose{}$. However, we motivated function-space VI to avoid weight-space priors, and in this case the bound in Eq. 2.3 can be looser. We will indeed see below that the bound becomes infinitely loose in practice, and we thus propose a different objective in Section 3.
@@
Two intractabilities prevent directly maximizing the ELBO in function space (Eq 2.3). First, it is not obvious how to practically evaluate or estimate the KL divergence between two measures in function space in Eq 2.4. Sun et al., (2019) showed that it can be expressed as a supremum of KL divergences between finite-dimensional distributions,
$D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\sup_{{\bm{x}}\in\mathcal{X}^{M}}D_{\mathrm{KL}}\mathopen{}\left(q_{\phi}(f({\bm{x}}))\,\|\,p(f({\bm{x}}))\right)\mathclose{}.$ (2.5)
@@
Here, ${\bm{x}}=\{{\bm{x}}^{(i)}\}_{i=1}^{M}\in\mathcal{X}^{M}$ is a set of $M$ points in feature space $\mathcal{X}$, and $q_{\phi}(f({\bm{x}}))$ and $p(f({\bm{x}}))$ are densities of the marginals of $\mathbb{Q}_{\phi}$ and $\mathbb{P}$ on $\{f({\bm{x}}^{(i)})\}_{i=1}^{M}$ respectively. To obtain a tractable approximation of the supremum over infinitely many sets, Sun et al., (2019) replace it by an expectation and Rudner et al., 2022b estimate it from samples.
@@
Second, we cannot express the pushforward measure $\mathbb{Q}_{\phi}$ of $q_{\phi}({\bm{w}})$ in a closed form because the neural network is nonlinear. Previous work has proposed to mitigate this issue using implicit score function estimators (Sun et al.,, 2019) or a linearized BNN $f_{L}$ to obtain a closed-form Gaussian variational measure (Rudner et al., 2022a, ; Rudner et al., 2022b, ). Our proposal in Section 3 follows the linearized BNN approach as it only minimally modifies the BNN, preserving most of its inductive bias (Maddox et al.,, 2021) while considerably simplifying the problem by turning the pushforward of $q_{\phi}({\bm{w}})$ into a Gaussian process (GP). More specifically, we consider a Gaussian variational distribution $q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}=\mathcal{N}\mathopen{}\left({\bm{m}},{\bm{S}}\right)\mathclose{}$ with parameters $\phi=\left\{{\bm{m}},{\bm{S}}\right\}$, and we obtain $f_{L}$ by linearizing $f$ in the weights around ${\bm{w}}={\bm{m}}$,
$f_{L}({\bm{x}};{\bm{w}}):=f({\bm{x}};{\bm{m}})+J({\bm{x}};{\bm{m}})({\bm{w}}-{\bm{m}})\qquad\text{where}\qquad J({\bm{x}};{\bm{m}})=\nabla_{\!{\bm{w}}}f({{\bm{x}};{\bm{w}}})|_{{\bm{w}}={\bm{m}}}.$ (2.6)
@@
Thus, for ${\bm{w}}\sim q_{\phi}({\bm{w}})$ we have, $f_{L}({\bm{x}};{\bm{w}})\sim\mathcal{N}\mathopen{}\left(f({\bm{x}};{\bm{m}}),J({\bm{x}};{\bm{m}}){\bm{S}}J({\bm{x}};{\bm{m}})^{\top}\right)\mathclose{}\;\forall{\bm{x}}$, and the function $f_{L}(\,\cdot\,;{\bm{w}})$ is a degenerate GP (as $\operatorname{rank}(J(\,\cdot\,;{\bm{m}}){\bm{S}}J(\,\cdot\,;{\bm{m}})^{\top})\leq\text{number\leavevmode\nobreak\ of\leavevmode\nobreak\ weights}<\infty$),
@@
$D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ is infinite in most practically relevant cases.
@@
Burt et al., (2020) point out an even more severe issue of function-space VI in BNNs: $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ is in fact infinite in many relevant cases, in particular for non-degenerate GP-priors with BNNs. The proof in Burt et al., (2020) is somewhat involved, but the fundamental reason for $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\infty$ is that $\mathbb{Q}_{\phi}$ has support on a finite-dimensional submanifold of the infinite-dimensional function space, while the measure $\mathbb{P}$ induced by a (non-degenerate) GP prior has support on the entire function space. That such a dimensionality mismatch can lead to infinite KL divergence can be seen in a finite-dimensional example: consider the KL-divergence between two Gaussians in $\mathbb{R}^{n}$ for some $n\geq 2$, one of which has support on the entire $\mathbb{R}^{n}$ (i.e., its covariance matrix ${\bm{\Sigma}}_{1}$ has full rank) while the other one has support only on a proper subspace of $\mathbb{R}^{n}$ (i.e., its covariance matrix ${\bm{\Sigma}}_{2}$ is singular). The KL divergence between two multivariate Gaussians has a closed form expression (Eq 2.9 with ${\gamma=0}$) that contains $\log\det\mathopen{}\left({\bm{\Sigma}}_{2}^{-1}{\bm{\Sigma}}_{1}\right)\mathclose{}$, which is infinite for singular ${\bm{\Sigma}}_{2}$.
@@
We find that the fact that $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\infty$ has severe practical consequences even when the KL divergence is only estimated from samples. It naturally explains the stability issues discussed in Appendix D.1 of Sun et al., (2019) (we compare the authors’ solution to this stability issue to our method in Section 3.2). Surprisingly, similar complications arise even in the setup by Rudner et al., 2022b , which performs VI in function space with the pushforward of a weight-space prior. While the KL divergence is technically finite because prior and variational posterior have the same support, numerical errors lead to mismatching supports and thus to stability issues even there.
@@
In summary, the ELBO for VI in BNNs is not well-defined for a large class of interesting function-space priors. In section 3, we propose a solution by using the so-called regularized KL divergence, which we introduce next.
2.2 Regularized KL divergence for Gaussian measures

### v3

# Well-Defined Function-Space Variational Inference in Bayesian Neural Networks via Regularized KL-Divergence

@@ Abstract
Bayesian neural networks (BNN) promise to combine the predictive performance of neural networks with principled uncertainty modeling crucial for safety-critical systems and decision making. However, posterior uncertainties depend on the choice of prior, and finding informative priors in weight-space has proven difficult. This has motivated variational inference (VI) methods that pose priors directly on the function represented by the BNN rather than on weights. In this paper, we address a fundamental issue with such function-space VI approaches pointed out by Wild et al. [2022a], who showed that the objective function (ELBO) is negative infinite for most priors of interest. Our solution builds on generalized VI with the regularized KL divergence and is, to the best of our knowledge, the first well-defined variational objective for inference in BNNs with Gaussian process (GP) priors. Experiments show that our method successfully incorporates the properties specified by the GP prior, and that it provides competitive uncertainty estimates for regression, classification and out-of-distribution detection compared to BNN baselines with both function and weight-space priors.
@@ 1 Introduction
Figure 1: Inference with our GFSVI on synthetic data (gray circles) with Gaussian process priors encoding different properties such as smoothness (increasing from Matérn-1/2 to RBF) and periodicity (last panel).
@@
Neural networks have shown impressive results in many fields but fail to provide well-calibrated uncertainty estimates, which are essential in applications associated with risk, such as healthcare [Abdullah et al., 2022] or finance [Bew et al., 2019]. Bayesian neural networks (BNNs) offer to combine the scalability and predictive performance of neural networks with principled uncertainty modeling by explicitly capturing epistemic uncertainty, which results from finite training data. While the choice of prior strongly affects posterior uncertainties, specifying informative priors on BNN weights has proven difficult and is hypothesized to have limited their practical applicability [Knoblauch et al., 2022, Tran et al., 2022]. For instance, the default isotropic Gaussian prior, which is often chosen for tractability rather than for the beliefs it carries [Knoblauch et al., 2022], is known to have pathological behavior in some cases [Cinquin et al., 2021, Tran et al., 2022]. A promising approach to solve this issue is to place priors directly on the function represented by the BNN instead of the weights. Function-space priors allow incorporating interpretable knowledge, for instance using the Gaussian Process (GP) literature to improve prior design and selection [Williams and Rasmussen, 2006].
@@
A recent line of work has focused on using function-space priors in BNNs with variational inference (VI) [Sun et al., 2019]. VI is appealing because of its successful application to BNNs, its flexibility in terms of approximate posterior parameterization, and its scalability to large datasets and models [Hoffman et al., 2013, Blundell et al., 2015]. Unfortunately, for BNNs with function-space priors, the Kullbach-Leibler (KL) divergence term in the VI objective (ELBO) involves two intractabilities: (i) a supremum over infinitely many subsets and (ii) access to the density of the distribution of the BNN’s function, which has no closed-form expression. Sun et al. [2019] propose to address problem (i) by approximating the supremum in the KL divergence by an expectation, and problem (ii) by using implicit score function estimators (which make this method difficult to use in practice [Ma and Hernández-Lobato, 2021]). However, the problem is actually more severe. Not only is the KL divergence intractable, it is infinite in most cases of interest [Wild et al., 2022a], such as when the prior is a non-degenerate GP or a BNN with a different architecture. Thus, in these (and many more) situations, the KL divergence cannot even be approximated. As a consequence, more recent work abandons using BNNs and instead uses deterministic neural networks to parameterize basis functions [Ma and Hernández-Lobato, 2021] or a GP mean [Wild et al., 2022b]. The only prior work [Rudner et al., 2022b] that overcomes the issue pointed out by Wild et al. [2022a] does so by deliberately limiting itself to cases where the KL divergence is known to be finite (by defining the prior as the pushforward of a weight-space distribution). Therefore, the method by Rudner et al. [2022b] suffers from the same issues regarding prior specification as other weight-space inference method.
@@
In this paper, we address the argument by Wild et al. [2022a] that VI does not provide a valid objective for inference in BNNs with genuine function-space priors, and we propose to apply the framework of generalized VI [Knoblauch et al., 2022]. We present a simple method for function-space inference with GP priors that builds on the regularized KL divergence [Quang, 2019], which generalizes the conventional KL divergence and is finite for any pair of Gaussian measures. We obtain a Gaussian measure for the variational posterior by considering the linearized BNN from Rudner et al. [2022b], and we are free to choose a function-space prior from a large set of GPs which have an associated Gaussian measure on the considered function space. While the regularized KL divergence is still intractable, it can be consistently estimated from samples with a known error bound. We find that our method effectively incorporates the beliefs specified by GP priors (see figure 1, discussed further in section 4) and that it yields competitive performance compared to BNN baselines. To the best of our knowledge, our method is the first to provide a well-defined objective for function-space inference in BNNs with informative GP priors. Our contributions are summarized below:
@@
    We use generalized VI with the regularized KL divergence to mitigate the issue of an infinite KL divergence when using VI in BNNs with function-space priors.
@@
    We present a new and well-defined objective for function-space inference in the linearized BNN with GP priors, resulting in a simple algorithm.
@@
    We show that our method accurately captures structural properties specified by the GP prior and provides competitive uncertainty estimates for regression, classification, and out-of-distribution detection compared to baselines with both function- and weight-space priors.
The paper is structured as follows: section 2 introduces function-space VI and the regularized KL divergence; section 3 presents our method for generalized function-space VI (GFSVI) in BNNs; section 4 reports experimental results; section 5 discusses related work; and section 6 concludes.
@@ 2 Background
In this section, we provide background on function-space variational inference in BNNs and discuss the fundamental issue of an infinite KL divergence. We then introduce the regularized KL divergence, which is the basis for our solution presented in Section 3.
@@ 2.1 Function-space VI in BNNs
We consider a neural network $f({\,\cdot\,;{\bm{w}}})$ with weights ${\bm{w}}\in{}^{p}$, and a data set $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ with features ${\bm{x}}_{i}\in\mathcal{X}\subset{}^{d}$ and values $y_{i}\in\mathcal{Y}$. Bayesian Neural Networks are specified further by a likelihood function $p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}=\prod_{i=1}^{N}p\mathopen{}\left(y_{i}\,|\,f({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}$ and—traditionally—a prior $p\mathopen{}\left({\bm{w}}\right)\mathclose{}$ on the weights, and one seeks the posterior distribution $p\mathopen{}\left({\bm{w}}\,|\,\mathcal{D}\right)\mathclose{}\propto p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}\,p\mathopen{}\left({\bm{w}}\right)\mathclose{}$. The method proposed in this paper builds on variational inference, which approximates $p\mathopen{}\left({\bm{w}}\,|\,\mathcal{D}\right)\mathclose{}$ with a variational distribution $q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}$, whose variational parameters $\phi$ maximize the evidence lower bound (ELBO),
@@
where $D_{\text{KL}}$ is the Kullback-Leibler divergence,
@@
Function-space variational inference.
Since neural network weights are not interpretable, we replace the weight-space prior $p({\bm{w}})$ with a prior $\mathbb{P}$ directly on the function $f\mathopen{}\left(\,\cdot\,;{\bm{w}}\right)\mathclose{}$, which we denote simply as $f$ when there is no ambiguity. Here, the symbol $\mathbb{P}$ denotes a probability measure that does not admit a density since the function space is infinite-dimensional. We compute the expected log-likelihood as in the first term of Eq. 2.1. For the KL-term (Eq. 2.2), a naive VI-method would use the pushforward of $q_{\phi}({\bm{w}})$ along ${\bm{w}}\mapsto f\mathopen{}\left(\,\cdot\,;{\bm{w}}\right)\mathclose{}$, which defines the variational measure $\mathbb{Q}_{\phi}$, resulting in the ELBO in function space,
@@
with $D_{\text{KL}}$ the KL divergence between measures
@@
Here, the Raydon-Nikodym derivative $d\mathbb{Q}_{\phi}/d\mathbb{P}$ generalizes the density ratio from Eq. 2.2. Like Eq. 2.1, the ELBO in Eq. 2.3 is a lower bound on the evidence [Wild et al., 2022a]. In fact, if $\mathbb{P}$ is the push-forward of $p({\bm{w}})$ then Eq. 2.3 is a tighter bound than Eq. 2.1 by the data processing inequality, $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}\leq D_{\mathrm{KL}}\mathopen{}\left(q_{\phi}\,\|\,p\right)\mathclose{}$. However, we motivated function-space VI to avoid weight-space priors, and in this case the bound in Eq. 2.3 can be looser. We will indeed see below that the bound becomes infinitely loose in practice, and we thus propose a different objective in Section 3.
@@
Two intractabilities prevent directly maximizing the ELBO in function space (Eq 2.3). First, it is not obvious how to evaluate or estimate the KL divergence between two measures in Eq 2.4. Sun et al. [2019] showed that it can be expressed as a supremum of KL divergences between finite-dimensional distributions,
$D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\sup_{{\bm{x}}\in\mathcal{X}^{M},M\in\mathbb{N}}D_{\mathrm{KL}}\mathopen{}\left(q_{\phi}(f({\bm{x}}))\,\|\,p(f({\bm{x}}))\right)\mathclose{}.$ (2.5)
@@
Here, ${\bm{x}}=\{{\bm{x}}^{(i)}\}_{i=1}^{M}\in\mathcal{X}^{M}$ is a set of $M$ points in feature space $\mathcal{X}$, and $q_{\phi}(f({\bm{x}}))$ and $p(f({\bm{x}}))$ are densities of the marginals of $\mathbb{Q}_{\phi}$ and $\mathbb{P}$ on $\{f({\bm{x}}^{(i)})\}_{i=1}^{M}$ respectively. Sun et al. [2019] approximates the supremum over infinitely many sets by an expectation, and Rudner et al. [2022b] estimates it from samples.
@@
Second, we cannot express the pushforward measure $\mathbb{Q}_{\phi}$ in closed form because the neural network is nonlinear. Previous work has proposed to mitigate this issue using implicit score function estimators [Sun et al., 2019] or a linearized BNN $f_{L}$ to obtain a closed-form Gaussian variational measure [Rudner et al., 2022a, b]. Our proposal in Section 3 follows the linearized BNN approach as it only minimally modifies the BNN, preserving most of its inductive bias [Maddox et al., 2021] while considerably simplifying the problem by turning the pushforward of $q_{\phi}({\bm{w}})$ into a GP. More specifically, we consider a Gaussian variational distribution $q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}=\mathcal{N}\mathopen{}\left({\bm{m}},{\bm{S}}\right)\mathclose{}$ with parameters $\phi=\left\{{\bm{m}},{\bm{S}}\right\}$, and we define a linearized BNN $f_{L}$ by linearizing $f$ as a function of the weights around ${\bm{w}}={\bm{m}}$,
$f_{L}({\bm{x}};{\bm{w}}):=f({\bm{x}};{\bm{m}})+J({\bm{x}};{\bm{m}})({\bm{w}}-{\bm{m}})$ (2.6)
@@
with $J({\bm{x}};{\bm{m}})=\nabla_{\!{\bm{w}}}f({{\bm{x}};{\bm{w}}})|_{{\bm{w}}={\bm{m}}}$. Thus, ${\bm{w}}\sim q_{\phi}({\bm{w}})$ implies $f_{L}({\bm{x}};{\bm{w}})\sim\mathcal{N}\mathopen{}\left(f({\bm{x}};{\bm{m}}),J({\bm{x}};{\bm{m}}){\bm{S}}J({\bm{x}};{\bm{m}})^{\top}\right)\mathclose{}$ for all ${\bm{x}}\in\mathcal{X}$, and so the function $f_{L}(\,\cdot\,;{\bm{w}})$ is a degenerate GP (as $\operatorname{rank}(J(\,\cdot\,;{\bm{m}}){\bm{S}}J(\,\cdot\,;{\bm{m}})^{\top})\leq p<\infty$),
@@
$D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ is infinite in most relevant cases.
@@
Wild et al. [2022a] point out an even more severe issue of function-space VI in BNNs: $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ (Eq. 2.4) is in fact infinite in most relevant cases, in particular for non-degenerate GP-priors. Thus, approximating $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ in these settings is futile. Their proof is somewhat involved, but the fundamental reason for $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\infty$ is that $\mathbb{Q}_{\phi}$ has support on a finite-dimensional submanifold of the infinite-dimensional function space, while the measure $\mathbb{P}$ induced by a (non-degenerate) GP prior has support on the entire function space. That such a dimensionality mismatch can lead to infinite KL divergence can already be seen in a finite-dimensional example: consider the KL-divergence between two Gaussians in $\mathbb{R}^{n}$ for $n\geq 2$, one of which has support on the entire $\mathbb{R}^{n}$ (i.e., its covariance matrix ${\bm{\Sigma}}_{1}$ has full rank) while the other one has support only on a proper subspace of $\mathbb{R}^{n}$ (i.e., its covariance matrix ${\bm{\Sigma}}_{2}$ is singular). The KL divergence between multivariate Gaussians has a closed form expression (Eq. 2.9 with ${\gamma=0}$) that contains $\log\det\mathopen{}\left({\bm{\Sigma}}_{2}^{-1}{\bm{\Sigma}}_{1}\right)\mathclose{}$, which is infinite for singular ${\bm{\Sigma}}_{2}$.
@@
We find that the fact that $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\infty$ has severe practical consequences even when the KL divergence is only estimated from finite samples. It naturally explains the stability issues discussed in Appendix D.1 of Sun et al. [2019] (we compare the authors’ solution to this stability issue to our method in Section 3.2). Surprisingly, similar complications arise even in the setup by Rudner et al. [2022b], which performs VI in function space with the pushforward of a weight-space prior. While this makes the KL divergence technically finite because prior and variational posterior have the same support, numerical errors lead to mismatching supports and thus to stability issues even there.
@@
In summary, the ELBO for VI in BNNs is not well-defined for most interesting function-space priors. In section 3, we propose a solution by using the so-called regularized KL divergence, which we introduce next.
2.2 Regularized KL divergence